OpenStudy (anonymous):

what is the derivative of y=x/(1+cos(x))

OpenStudy (anonymous):

\[y=\frac{ x }{ 1+\cos x }\] \[\frac{ dy }{ dx }=\frac{ \left( 1+\cos x \right)*1-x*(-\sin x) }{ \left( 1+\cos x \right)^2 }=\frac{ 1+\cos x+x \sin x }{ \left( 1+\cos x \right)^2 }\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw